ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 159, 2007

Abraham, Uri and Rene Schipperus Infinite games on finite sets  205

Avilés, Antonio The number of weakly compact sets which generate a Banach space

189

Bernig, Andreas The normal cycle of a compact definable set  373

Frauenfelder, Urs and Felix Schlenk Hamiltonian dynamics on convex symplectic manifolds

1

Howard, R. and D. Hug Smooth convex bodies with proportional projection functions

317

Hug, D. See R. Howard

Joswig, Michael and Thilo Rörig Neighborly cubical polytopes and spheres

221

Lee, Han Ju Complex convexity and monotonicity in quasi-Banach lattices

57

Lerner, Andrei K. BMO-boundedness of the maximal operator for arbitrary measures

243

Loke, Hung Yean and Gordan Savin On local lifts from G_{2}(ℝ) to

Sp_{6}(ℝ) and F_{4}(ℝ) 349

Loukaki, Maria On distinct character degrees  93
Mattarei, Sandro Inverse-closed additive subgroups of fields  343
Mitra, Sudeb Extensions of holomorphic motions  277

Rörig, Thilo See Michael Joswig

Rosłanowski, A. and S. Shelah Sheva-Sheva-Sheva: Large creatures 109

Rovenski, Vladimir Submanifolds with restrictions on extrinsic q-th scalar curvature

289

Savin, G. See Hung Yean Loke

Schipperus, Rene See Uri Abraham

Schlenk, Felix See Urs Frauenfelder

Schraudner, M. The canonical-boundary representation for automorphism groups of locally compact countable state Markov shifts

253

Shelah, S. See A. Rosłanowski

Wang, Y. See H. Wei

Wei, H. and Y. Wang On CAS-subgroups of finite groups  175